OpenStudy (anonymous):

how many liters each of 15% acid and 33% acid should be mixed to get 120 liters of 21% acid? I get x=80 and y is equal to 40 is this correct?

OpenStudy (anonymous):

0.15*x+.33*y = .21(x+y) x amountof 15% + y amount of 33% = (x+y) abount of 21%

OpenStudy (anonymous):

x+y=120

OpenStudy (anonymous):

and is the second one supposed to look like 0.15x+0.33y=0.21?